OpenStudy (ilovemath14):

Find the midpoint 1. (2,3) (5,8) 2. (-7,5) (-2,-5) 3. (2,-4) (2,5)

OpenStudy (matthew120):

is there a pic

Mehek (mehek14):

Add the x values and y values and divide by 2 Start with 1 The x values are 2 and 5. Y values are 3 and 8 \((\dfrac{2+5}{2},\dfrac{3+8}{2})\)

OpenStudy (ilovemath14):

No just that @matthew120

OpenStudy (ilovemath14):

So (3.5,5.5)? @mehek14

Mehek (mehek14):

Yep

Mehek (mehek14):

Second one: \((\dfrac{-7+-2}{2},\dfrac{5+-5}{2})\)

OpenStudy (ilovemath14):

(-4.5,2.5)?

Mehek (mehek14):

5 + -5 = 0 0/2 = 0 But -4.5 is right

Mehek (mehek14):

So (-4.5, 0)

OpenStudy (ilovemath14):

oh my bad @mehek14

Mehek (mehek14):

Last one Want to try on your own and lets see how it turns out? :)

OpenStudy (ilovemath14):

Sure 2+2 5+-4 ---- ----- 2 , 2 4 1 --- --- 2 , 2 Answer: (2, 0.5) Am I correct?

OpenStudy (ilovemath14):

@mehek14

Mehek (mehek14):

Right (: